OpenStudy (anonymous):

find the radius of the convergence of the Taylor series around x=0 for ln(1-x)

OpenStudy (anonymous):

i guess it cannot be bigger than one, since log (0) i undefined.

OpenStudy (anonymous):

we can compute it if you like

OpenStudy (anonymous):

standard trick is to take the derivative get \[-\frac{1}{1-x}\], write the expansion for that one and then integrate term by term

OpenStudy (anonymous):

get \[-\sum\frac{x^k}{k}\]

OpenStudy (anonymous):

ratio test immediately tells you radius of convergence is 1